OpenStudy (anonymous):

What is the square root of 93?

OpenStudy (anonymous):

if you want approximate value, then \[\sqrt{93} = 9.643650761...\] 93 = 3*31, 31 is simple, so there is no chanses to simplify this, i.e. \[\sqrt{93}=\sqrt{3}*\sqrt{31}\]